Problem 3-2A (Video) (Part Level Submission)

Rosenthal Company manufactures bowling balls through two processes: Molding and Packaging. In the Molding Department, the urethane, rubber, plastics, and other materials are molded into bowling balls. In the Packaging Department, the balls are placed in cartons and sent to the finished goods warehouse. All materials are entered at the beginning of each process. Labor and manufacturing overhead are incurred uniformly throughout each process. Production and cost data for the Molding Department during June 2020 are presented below.

Production Data

June

Beginning work in process units 0
Units started into production 29,480
Ending work in process units 2,680
Percent complete—ending inventory 40 %

Cost Data

Materials $265,320
Labor 71,824
Overhead 151,152
    Total $488,296

d) Determine the costs to be assigned to the units transferred out and in process for June

e) Prepare a production cost report for the Molding Department for the month of June
Answer #1
d
Transferred out 455600
Work in process, June 30 32696
e
ROSENTHAL COMPANY
Molding Department
Production Cost Report
For the Month Ended June 30, 2020
Equivalent Units
Physical Units Materials Conversion Costs
Units to be accounted for
Work in process, June 1 0
Started into production 29480
Total units 29480
Units accounted for
Transferred out 26800 26800 26800
Work in process, June 30 2680 2680 1072
Total units 29480 29480 27872
Costs Materials Conversion Costs Total
Unit costs:
Total costs 265320 222976 488296
Equivalent units 29480 27872
Unit costs 9.00 8.00 17.00
Costs to be accounted for
Work in process, June 1 0
Started into production 488296
Total costs 488296
Cost Reconciliation Schedule
Costs accounted for
Transferred out 455600 =26800*17
Work in process, June 30
      Materials 24120 =2680*9
      Conversion costs 8576 32696 =1072*8
Total costs 488296